OpenStudy (kingj):

What is the remainder when (2x3 – 3x2 – 10x + 15) ÷ (x + 2) ?

OpenStudy (anonymous):

use the remainder theorem: plug in x = -2 in the function - the result is the remainder

OpenStudy (anonymous):

remainder theorem states that when f (x) is divided by (x - a) the remainder is f(a).

OpenStudy (kingj):

Thank you but I need the answer just incase i mess up

OpenStudy (anonymous):

so here it is 2*(-2)^3 - 3(-2)^2 -10(-2) - 15 = remainder

OpenStudy (kingj):

oh ok

OpenStudy (anonymous):

= -16 -12 + 20 - 15 = -23

OpenStudy (anonymous):

sorry its + 15 so thas comes to -28 + 35 = 7

OpenStudy (anonymous):

ok?